OpenStudy (anonymous):

Find the values of the six trig functions of θ with the given constraint. tan θ is undefined. π less than θ less than 2π

OpenStudy (anonymous):

theta equals pi/2

OpenStudy (anonymous):

i think you can get the rest

OpenStudy (anonymous):

no i cant

OpenStudy (amistre64):

its the bottom half of a circle, tan is the slope of the angle and that is undefined when it is pointing straight up or straight down. 3pi/2 is the only undefined tan in the bottom of the circle, also known as 270 degrees

OpenStudy (amistre64):

sin = -1 cos = 0 tan = not gonna say sec = 1/cos = not defined csc = 1/sin = -1

OpenStudy (amistre64):

cot = 1/tan = 0